OpenStudy (anonymous):

linear congruence: complete soln: 4x(is equivalent to) 5(mod 21) @Mathematics

OpenStudy (anonymous):

when I did the EEA I got 4(-5) + 21(1) = 1

OpenStudy (anonymous):

but then you have to mult by 5

OpenStudy (anonymous):

then x=xnot+bn/d

OpenStudy (anonymous):

when I multiplied by 5 I got 4(-25) + 21(5) = 5

OpenStudy (anonymous):

o ya then xnot is -25

OpenStudy (anonymous):

x=-25+21n

OpenStudy (anonymous):

your way of writing things is different then how I did it

OpenStudy (anonymous):

you have to find all solutions

OpenStudy (anonymous):

yeah I got x is congruent to 17 (mod 21)

OpenStudy (anonymous):

k so where do i go from here

OpenStudy (anonymous):

from the EEA you just convert it back into congruent form I think

OpenStudy (anonymous):

so x is congruent to -25 is congruent to what

OpenStudy (anonymous):

so 4(-25) congruent to 5 (mod 21) so x congruent to -25 (mod 21/1) is the complete solution and that is equivalent to x congruent to 17 (mod21/1

OpenStudy (anonymous):

since (-25) is some solution for x

OpenStudy (anonymous):

then you just put it into the Linear Congruence Theorem, V1

OpenStudy (anonymous):

to get the complete solution

OpenStudy (anonymous):

o crap ok now i see it

OpenStudy (anonymous):

I don't know this stuff super well either lol

OpenStudy (anonymous):

did you show eea?

OpenStudy (anonymous):

yeah I put it in there

OpenStudy (anonymous):

but theres no room

OpenStudy (anonymous):

the EEA for those were really short so I just wrote it in there tiny

OpenStudy (anonymous):

have you done any other problems on the assignment? I'm stuck on 5c lol

OpenStudy (anonymous):

only up to 3

OpenStudy (anonymous):

ok well I guess I might try asking on here again then

OpenStudy (anonymous):

metoo

OpenStudy (anonymous):

4x(is equivalent to) 5(mod 21) 21 | 4x-5 21|4*(-4)-5 Hence -4 is a solution. ??? I am not sure.

OpenStudy (anonymous):

i dunno